OpenStudy (anonymous):

Solve 2x + 3y = 18 for y

OpenStudy (anonymous):

3y = 18 - 2x....Subtract 2x from both sides y = 1/3(18 - 2x)....Divide both sides by 3 y = 6 - 2x/3...Multiply through by 1/3

OpenStudy (saifoo.khan):

3y = 18 -2x y = 18 - 2x ------ 3